2(x-6)-4(x+2)=-8 solve for x
Answer by RedemptiveMath(80)

There is multiple ways you can do this problem. Using the distributive property:

2(x-6)-4(x+2)=-8
2x-12-4x-8=-8
-2x-20=-8
-2x=12
x=-6

Using division:

2(x-6)-4(x+2)=-8
{[2(x-6)]/2}-{[4(x+2)]/2}=-8/2
x-6-2(x+2)=-4
x-6-2x-4=-4
-x=6
x=-6